The energy difference between the ground state and first excited state of an
electron in a one-dimensional rigid box of length 10-8 cm is

Grade:

6 Answers

Arun Kumar IIT Delhi
9 years ago
Energy in one dimensional box is determined by
h is planck's constant
L is length of the box.
m is mass of electron

Arun Kumar
IIT Delhi
Askiitians Faculty
Mahesh
3 years ago
Calculate the energy difference between ground sstate and first excited state is an electron in box length 10power minus 8cm
 
Abhishek
2 years ago
En=n²h²/8mL²
En=n²×(6.63×10^-34)²/8×9.1×10^-34×(10^-8)²
En=3.75×10^-3×n²   eV
For ground state (n=1), E1=3.75×10^-3  eV
For exited state (n=2), E2=0.015 eV
E2-E1= 11.25  eV
Abhishek Kumbhar
one year ago
The duration of a radar pulse is 10-0 sec. Calculate tpe uncertainty in it's energy 
in electron volts. Given that Planck's constant h = 6.63 x 10·34 JS
Abhishek Kumbhar
one year ago
The duration of a radar pulse is 10-0 sec. Calculate tpe uncertainty in it's energy
in electron volts. Given that Planck's constant h = 6.63 x 10·34 JSec.
Abhishek Kumbhar
one year ago
The duration of a radar pulse is 10-0 sec. Calculate tpe uncertainty in it's energy 
in electron volts. Given that Planck's constant h = 6.63 x 10·34 JSec.